# AV-Master: Dual-Path Comprehensive Perception Makes Better Audio-Visual Question Answering

### v1

@@ Abstract
Audio-Visual Question Answering (AVQA) requires models to effectively utilize both visual and auditory modalities to answer complex and diverse questions about audio-visual scenes. However, existing methods lack sufficient flexibility and dynamic adaptability in temporal sampling and modality preference awareness, making it difficult to focus on key information based on the question. This limits their reasoning capability in complex scenarios. To address these challenges, we propose a novel framework named AV-Master. It enhances the model’s ability to extract key information from complex audio-visual scenes with substantial redundant content by dynamically modeling both temporal and modality dimensions. In the temporal dimension, we introduce a dynamic adaptive focus sampling mechanism that progressively focuses on audio-visual segments most relevant to the question, effectively mitigating redundancy and segment fragmentation in traditional sampling methods. In the modality dimension, we propose a preference-aware strategy that models each modality’s contribution independently, enabling selective activation of critical features. Furthermore, we introduce a dual-path contrastive loss to reinforce consistency and complementarity across temporal and modality dimensions, guiding the model to learn question-specific cross-modal collaborative representations. Experiments on four large-scale benchmarks show that AV-Master significantly outperforms existing methods, especially in complex reasoning tasks.
@@ I Introduction
Humans perceive the world through various modalities, such as vision, hearing, and touch. Inspired by such multisensory experiences, researchers have increasingly focused on a range of multimodal understanding tasks [1, 2, 3]. Among these tasks, Audio-Visual Question Answering (AVQA) [4, 5, 6] is a practical task with broad application prospects. AVQA utilizes both visual and auditory modalities and requires the model to discover the associations between them to answer various questions. This process involves dynamically understanding audio-visual segments and addressing question-specific modality preferences, significantly increasing the complexity of the task.
@@
For AVQA, we argue it is essential to focus on the following: (i) How can the model identify the most relevant visual and audio segments of the video related to the given question? (ii) How can the model extract cues from the most relevant modality information related to the given question? For (i), most current models employ discrete [7, 5] or continuous [6] sampling methods. Although these methods have achieved promising results, they still have notable limitations. The former compromises the inherent temporal cues in audio-visual segments, resulting in hallucinations during fine-grained understanding. The latter partially alleviates this issue but introduces significant redundancy, which hinders further enhancement of model performance. For (ii), most mainstream methods are not specifically designed to differentiate between various input modalities but instead treat them merely as supplementary information. A small number of researchers [8] have recognized these challenges and opted to design fusion weights for features from different modalities during the multimodal interaction process, thereby biasing the model towards a particular modality. However, this method overlooks the challenges in learning significant differences among fine-grained features and requires the model to reinterpret modality preferences in complex multimodal semantics, without explicitly enhancing the learning process, thus providing limited guidance for improving the model’s performance.
@@
To solve these issues, we propose AV-Master, which can find the important fine-grained areas related to the current question within complex dynamic audio-visual scenes, and generate the optimal answer by integrating global preference information. As shown in Fig. 1, unlike previous work [5, 6], AV-Master employs dynamic adaptive focus sampling to extract the fine-grained focus features from the audio-visual segments relevant to the current questioning scenario. The focus sampling combines the advantages of discrete (, uniform [7], top-k [5]) and continuous (, gaussian, gaussian-experts [6]) sampling methods, allowing the model to capture all continuous time steps while significantly reducing redundant information obtained from sampling. Furthermore, we propose a global preference activation strategy focused on model modality preferences. This strategy determines the preference distribution of the model in different modalities through independent perception and improves the decision-making capabilities of the model. Compared to using dynamic weights at the multimodal fusion stage, our enhancement strategy focuses on the global information in the initial features. This avoids biases caused by decoding from complex multimodal features, while also complementing the fine-grained information obtained through temporal dynamic perception, achieving comprehensive audio-visual understanding. Our contributions are summarized as follows:
@@ II-A Audio-Visual Scene Understanding
In recent years, audio-visual scene understanding has emerged as a significant research area, garnering ever-increasing attention from the academic community. The core idea of this field originates from the human instinct to perceive the world through the synergy of multiple senses. Specifically, visual and auditory information are not only complementary but also often inseparable in understanding complex, dynamic environments. These two modalities are tightly linked through semantic consistency (e.g., the image of a dog matching the sound of its bark) and spatio-temporal correlation (a sound synchronizing with and emanating from its visual source). This interplay provides the key elements for achieving a sophisticated understanding of scenes that surpasses the capabilities of single-modality perception. Building on this foundation, researchers have explored numerous specific sub-tasks such as sound source localization [9, 10, 11], action recognition [12, 13], event detection [14, 15, 16], video parsing [17, 18, 19], audio-visual source separation [20, 21], etc. These tasks aim to fully leverage the interaction and fusion of audio and visual information to overcome the limitations of unimodal perception, thereby enhancing the fine-grained understanding of dynamic audio-visual scenes.
Within this broad research landscape, our work focuses on the more advanced cognitive task of audio-visual question answering. We achieve precise scene understanding and reasoning based on multimodal information by specifically emphasizing the temporal dynamic perception of audio-visual segments and employing a global preference activation strategy to dynamically capture the dependency of different questions on specific modalities.
@@ II-B Audio-Visual Question Answering
The audio-visual question answering (AVQA) aims at achieving fine-grained comprehension and reasoning of complex audio-visual scenes. This task involves a comprehensive analysis of audio, visual, and their fused information, allowing the model to provide accurate answers to different questions. Existing research focuses mainly on unimodal audio question answering (AQA) and visual question answering (VQA), but a single modality is insufficient to fully capture the rich semantic information contained in natural videos. To this end, recent works [5, 8, 6] have begun exploring the field of AVQA. Among these works, Li et al. constructed declarative sentence prompts based on question templates to help the temporal awareness module better identify key segments relevant to the question, and designed a novel spatial awareness module to facilitate efficient fusion of visual tokens. Zhao et al. aligned audio-visual cues across spatial and temporal dimensions through contrastive learning, and adaptively assigned fusion weights to the visual and audio modalities according to the question. Kim et al. proposed the QA-TIGER framework, which adaptively focuses on both continuous and discontinuous frames according to the question, explicitly injects question information, and applies progressive optimization.
However, existing works only adopt simple audio-visual segment selection methods, which introduce a large amount of redundant information, making it difficult for the subsequent decoder to extract key clues from the coarse-grained multimodal features. Moreover, they overlook the fact that different questions may have varying demands on visual and audio modalities. As a result, the models often fail to dynamically leverage the relevant dominant modality for different questions and instead perform a coarse multimodal fusion, which, to some extent, undermines the overall performance of the model. In contrast, our work enables fine-grained key audio-visual learning through dynamic adaptive sampling based on the question, and enhances the model from a global perspective by leveraging modality preference, ultimately leading to more accurate answers.
@@ III-A Input Representation
where the performs summation along the sequence dimension, while the takes the maximum value along the feature dimension. The fused feature and represents the intermediate outputs.
@@ III-D Optimization and Answer Prediction
During training, the fused feature and the sentence-level question feature are jointly fed into the multimodal decoder to generate the answer prediction distribution, which is then used together with the ground-truth labels to compute the answer loss (denoted as ). In the preference activation path, features and are separately input into two independent audio/visual decoders, which also generate prediction distributions under the guidance of , resulting in two preference losses (denoted as and ). Moreover, a contrastive loss (denoted as ) is applied between the dynamic perception and preference activation paths, aiming to enhance the stability of the dual-path paradigm and improve the model’s discriminative ability by leveraging hard negative samples. During inference, we sum all the predicted distributions from the model and use the argmax function to obtain the final answer.
@@ IV-A Experimental Setting
(a) Dataset and evaluation metric: In this paper, we validate our proposal on four datasets: MUSIC-AVQA [7], MUSIC-AVQA-R [24], MUSIC-AVQA-v2.0 [25], and AVQA [26]. The details are summarized in Table I. Consistent with previous work, the ablation and analysis experiments are conducted on the MUSIC-AVQA dataset by default.
MUSIC-AVQA [7] is a large-scale AVQA dataset focused on multimodal understanding and reasoning in music performance scenarios. It contains 45,601 question-answer pairs distributed across 9,288 videos, with a total duration of over 150 hours. The videos are primarily collected from YouTube and cover 22 types of musical instruments (such as guitar, cello, marimba, etc.). Annotations are manually created by human annotators. The question-answer pairs are divided into three modality scenarios (audio, visual, and audio-visual), encompassing nine question types (such as existential, location, counting, etc.) and 33 question templates. The answer set includes 42 different answers. The questions require fine-grained scene understanding and spatiotemporal reasoning of both audio and visual content, for example: “Is a certain instrument present?” or “Where is the source of the sound located?”. The MUSIC-AVQA dataset emphasizes the interaction between audio and visual modalities in music performance scenes. Compared to other video question answering datasets (such as general video datasets), it places a stronger focus on audio-visual correlation.
MUSIC-AVQA-R [24] is an extended version of the MUSIC-AVQA, designed to evaluate the robustness of AVQA models and to address the issues of limited question expression variety and potential biases in the original dataset. It is constructed by rephrasing and splitting the questions in the MUSIC-AVQA test set. Specifically, researchers used AI tools to rephrase the 9,129 questions in the test set 25 times, generating semantically equivalent but more diverse expressions. Then, three annotators independently voted to retain most of the questions. As a result, the number of test questions expanded from 9,129 to 211,572, and the vocabulary size increased from 93 to 465. All questions are divided into two subsets: frequent (head) and rare (tail), classified based on the frequency of answer occurrences. Compared to MUSIC-AVQA, MUSIC-AVQA-R enhances question diversity and is better suited for testing AVQA models under conditions of diverse questions and imbalanced answer distributions.
MUSIC-AVQA-v2.0 [25] is an improved version of the MUSIC-AVQA, aimed at addressing the data bias issues in the original dataset and constructing a more balanced and challenging dataset. Researchers manually collected 1,230 new instrument performance videos (sourced from YouTube) and created 8,100 new question-answer pairs to supplement the original MUSIC-AVQA dataset. They ensured a more even answer distribution across each question category and subcategory, especially for binary questions, where the answers are nearly evenly distributed to avoid significant bias. Some of the videos were horizontally flipped to generate symmetric question-answer pairs, enhancing data diversity. MUSIC-AVQA-v2.0 resolves the answer bias problems present in the MUSIC-AVQA dataset, making it a more reliable benchmark suitable for testing model performance in unbiased AVQA tasks.
To evaluate the performance of AV-Master in scenarios beyond musical performance, we additionally introduce the AVQA dataset. AVQA [26] is a large-scale benchmark specifically designed for audio-visual question answering in real-world settings. The dataset contains 57,015 real-life videos and 57,335 question-answer pairs, with a total duration of over 158 hours. The videos are sourced from the VGG-Sound dataset [27] and cover 165 categories of daily activities and natural sounds. The questions are manually designed to ensure reliance on both modalities for reasoning, involving various types of relations such as existence, location, temporal, causal, and intentional. With its large scale, diversity, and high-quality human annotations, the dataset has become a widely used benchmark for evaluating multimodal fusion methods in complex real-world scenarios.
@@
(b) Implementation details: For a fair comparison, we follow previous work [6] and adopt a similar setup: videos are uniformly sampled at a rate of 1 frame per second. Audio representations are extracted using the pre-trained VGGish model [23], while visual inputs and corresponding questions are encoded through the CLIP-ViT-L/14 model [22]. All extracted features are projected into a 512-dimensional space via a linear transformation. The predefined templates, and , are both composed of a set of learnable embeddings, each with a length of 8 and randomly initialized. The model is optimized using Adam [28] and starts with a learning rate of 1e-4, which is reduced by a factor of 0.1 every 8 epochs. The batch size is set to 32, and the model is trained for 30 epochs. Our proposed model is trained on NVIDIA GeForce RTX 4090 and implemented in PyTorch.
@@ IV-B Quantitative Results
# Methods Average Accuracy (%) A-QA V-QA AV-QA Avg 1 w/o. ALL 73.37 79.23 69.82 72.94 2 w/o. AVFC 79.83 85.55 74.04 78.11 3 w/o. DPCL 78.83 85.71 73.78 77.84 4 w/o. GPAP 77.90 85.05 73.10 77.12 5 w/o. TDPP 78.77 83.73 72.94 76.83 6 AV-Master 79.95 86.58 74.22 78.51
@@
(c) MUSIC-AVQA-R: As shown in Table III, compared to the MUSIC-AVQA dataset, AV-Master demonstrates more significant improvements (+3.20%) on the MUSIC-AVQA-R dataset, achieving an overall accuracy of 71.19% and setting a new state-of-the-art performance. The notable performance gain can be attributed to AV-Master’s dual-path learning paradigm, which provides the model with powerful generalization capabilities.
@@
As shown in Table VI, AV-Master achieves competitive performance (78.5% accuracy) compared to various pretraining-based models, despite having significantly fewer parameters (61M) and not relying on large-scale pretraining. In contrast, methods such as CAT+ and AVicuna, which are based on large language models, achieve notably lower performance (50.1% and 49.6% accuracy) under zero-shot conditions. Although these approaches leverage abundant resources and powerful backbone models, they often exhibit weaker generalization capabilities in domain-specific audio-visual tasks. Other pretraining-based models achieve impressive performance (80.7% accuracy) after fine-tuning on the MUSIC-AVQA dataset. However, our method achieves comparable results using significantly less training data, fewer trainable parameters, and a more lightweight audio-visual feature encoder. This highlights the practicality of AV-Master and its suitability for resource-constrained audio-visual scenarios.
# Lengths Average Accuracy (%) A-QA V-QA AV-QA Avg 1 16 80.51 85.96 74.04 78.34 2 12 80.26 85.88 74.08 78.30 3 8 79.95 86.58 74.22 78.51 4 4 80.63 85.22 73.92 78.10 5 2 80.38 85.84 73.98 78.26

### v2

@@ Abstract
Audio-Visual Question Answering (AVQA) requires models to effectively utilize both visual and auditory modalities to answer complex and diverse questions about audio-visual scenes. However, existing AVQA methods still face two fundamental challenges: accurately locating question-relevant temporal evidence in long audio-visual streams and adaptively exploiting the dominant modality when audio and visual cues contribute unequally across different questions. This limits their reasoning capability in complex scenarios. To address these challenges, we propose a novel framework named AV-Master. It enhances the model’s ability to extract key information from complex audio-visual scenes with substantial redundant content by dynamically modeling both temporal and modality dimensions. In the temporal dimension, we introduce a dynamic adaptive focus sampling mechanism that progressively focuses on audio-visual segments most relevant to the question, effectively mitigating redundancy and segment fragmentation in traditional sampling methods. In the modality dimension, we propose a preference-aware strategy that models each modality’s contribution independently, enabling selective activation of critical features. Furthermore, we introduce a dual-path contrastive loss to enhance the collaborative consistency between the temporal dynamic perception path and the global preference activation path, while simultaneously preserving the complementary audio and visual cues for answer prediction. Experiments on four large-scale benchmarks show that AV-Master significantly outperforms existing methods, especially in complex reasoning tasks. The code is available at: https://github.com/AoKoo233/AV-Master.
@@ I Introduction
Humans perceive the world through various modalities, such as vision, hearing, and touch. Inspired by such multisensory experiences, researchers have increasingly focused on a range of multimodal understanding tasks [64, 44, 32]. Among these tasks, Audio-Visual Question Answering (AVQA) [25, 27, 19] is a practical task with broad application prospects. AVQA leverages both visual and auditory modalities, requiring models to perform fine-grained, question-conditioned discrimination over short temporal segments and subtle modality-specific cues, rather than coarse understanding of entire video clips as in other tasks (e.g., video captioning [65]). For instance, to answer a question such as “Which flute produces sound first?”, the model must identify a specific brief moment and distinguish subtle cross-modal differences among similar candidate objects. This process involves dynamically understanding audio-visual segments and addressing question-specific modality preferences, significantly increasing the complexity of the task.
@@
For AVQA, we argue it is essential to focus on the following: (i) How can the model identify the most relevant visual and audio segments of the video related to the given question? (ii) How can the model appropriately model the question-conditioned modality dependency, thereby flexibly utilizing audio-specific, visual-specific, and joint multimodal evidence for answer prediction? For (i), most current models employ discrete [26, 24] or continuous [19] sampling methods. Although these methods have achieved promising results, they still have notable limitations. The former compromises the inherent temporal cues in audio-visual segments, resulting in hallucinations during fine-grained understanding. The latter partially alleviates this issue but introduces significant redundancy, which hinders further enhancement of model performance. For (ii), most mainstream methods are not specifically designed to differentiate between various input modalities but instead treat them merely as supplementary information. A few studies [67] have recognized these challenges and design modality-specific fusion weights during multimodal interaction, thereby guiding the model to favor a particular modality. However, such methods typically estimate modality importance only after multimodal interaction has been completed. At this stage, audio and visual evidence are already intertwined, so the resulting weights mainly serve to rebalance the fused features rather than explicitly enhancing modality-specific evidence that is closely related to the given question. As a result, their ability to guide performance improvement is limited.
@@
To solve these issues, we propose AV-Master, which can find the important fine-grained areas related to the current question within complex dynamic audio-visual scenes, and generate the optimal answer by integrating global preference information. As shown in Fig. 1, unlike previous work [24, 19], AV-Master employs dynamic adaptive focus sampling to extract the fine-grained focus features from the audio-visual segments relevant to the current questioning scenario. The focus sampling combines the advantages of discrete (, uniform [26], top-k [24]) and continuous (, gaussian, gaussian-experts [19]) sampling methods, allowing the model to capture all continuous time steps while significantly reducing redundant information obtained from sampling. Furthermore, we propose a global preference activation strategy focused on model modality preferences. This strategy determines the preference distribution of the model in different modalities through independent perception and improves the decision-making capabilities of the model. Compared to using dynamic weights at the multimodal fusion stage, our enhancement strategy focuses on the global information in the initial features. This design enables the model to learn audio and visual preferences under question guidance based on their respective unimodal representations, thereby providing complementary global cues to the temporal dynamic perception path that focuses on capturing fine grained temporal evidence. Our contributions are summarized as follows:
@@ II-A Audio-Visual Scene Understanding
Audio-visual scene understanding aims to learn semantically consistent and temporally aligned representations from visual and auditory streams. Early studies in this area mainly focused on cross-modal correspondence and low-level perception, such as sound source localization and audio-visual synchronization, where the goal is to associate a sound with its visual origin or to determine whether the two modalities are temporally matched [40, 9, 12]. Building on this line of research, subsequent works extended audio-visual learning to more structured perception tasks, including action recognition [13, 4, 42], event detection [46, 50, 70], video parsing [48, 21, 69], and audio-visual source separation [10, 11, 3]. More recent efforts further emphasize robust multimodal representation learning in complex dynamic environments, promoting the transition from simple cross-modal alignment to higher-level scene understanding [55, 59, 62]. This evolution shows that audio-visual learning has gradually moved from low-level correspondence modeling toward semantic reasoning over dynamic scenes. However, most existing audio-visual scene understanding tasks are still formulated as recognition, localization, or correspondence problems. In contrast, audio-visual question answering requires the model not only to align audio and visual cues, but also to identify question-relevant evidence, reason over temporally distributed events, and flexibly determine which modality should dominate the final decision. Therefore, AVQA can be viewed as a more challenging extension of audio-visual scene understanding.
@@ II-B Audio-Visual Question Answering
Audio-Visual Question Answering (AVQA) extends traditional visual question answering [14] (VQA) and audio question answering [23] (AQA) by requiring joint reasoning over video, audio, and language. Early attempts largely followed the paradigm of multimodal fusion developed for VQA or video reasoning, where models aggregate audio-visual features and answer questions based on global fused representations. Representative methods such as HME [7], PSAC [29] and HCRN [22] demonstrated the feasibility of multimodal question answering, but their reasoning process was still dominated by relatively coarse feature aggregation and lacked AVQA-specific temporal modeling.
With the release of dedicated benchmarks, research shifted from general multimodal fusion to task-specific spatiotemporal reasoning. ST-AVQA [26] explicitly introduced spatiotemporal modeling for music-oriented AVQA, showing that fine-grained temporal and spatial cues are essential for answering audio-visual questions. LAVISH [33] further strengthened audio-visual-language interaction, improving multimodal fusion for more challenging reasoning scenarios. These works established the importance of structured multimodal reasoning, but they still mainly relied on holistic feature encoding and fusion.
More recent AVQA methods have increasingly focused on identifying question-relevant temporal evidence. TSPM [24] improves AVQA by constructing template-based prompts and enhancing temporal/spatial perception, thereby helping the model locate informative segments related to the question. QA-TIGER [19] further explores adaptive temporal selection by combining continuous and discontinuous frame modeling, which alleviates part of the redundancy problem in dense temporal reasoning. Along another direction, AVAF-Net [67] introduces adaptive fusion weights to account for the varying importance of audio and visual modalities under different questions. Recent methods such as SHMamba [54], PSOT [31], and CoQo [38] continue this trend by enhancing sequence modeling, training strategy, or backbone representation, leading to stronger overall AVQA performance.
Despite this progress, two limitations remain insufficiently addressed. First, existing methods often model temporal evidence either through coarse discrete selection or dense continuous aggregation. The former may destroy temporal continuity and miss subtle yet important cues, while the latter tends to introduce considerable redundancy, making subsequent reasoning more difficult. Second, although some methods consider modality importance during multimodal fusion, modality preference is usually inferred only after the two modalities have been entangled. As a result, the model may still struggle to explicitly capture question-conditioned modality dominance from the original unimodal representations. Different from previous methods, our approach addresses these two issues in a unified manner by combining dynamic adaptive focus sampling for fine-grained temporal perception with a global preference activation strategy for modality-aware reasoning.
@@ III-A Input Representation
where the performs summation along the sequence dimension, while the takes the maximum value along the feature dimension. The is a broadcasting step along the sequence dimension and is used only for shape alignment before element wise addition. The fused feature and represents the intermediate outputs.
@@ III-C Global Preference Activation Path
Input:
Visual segment features ;
Audio segment features ;
Word-level question features ;
Learnable temporal bias .
Predefined templates ;
Parameter:
The lengths of predefined audio-visual templates
Output:
Fused collaborative audio-visual feature .
@@ III-D Optimization and Answer Prediction
During training, the fused feature and the sentence-level question feature are jointly fed into the multimodal decoder to generate the answer prediction distribution, which is then used together with the ground-truth labels to compute the answer loss (denoted as ). In the preference activation path, features and are separately input into two independent audio/visual decoders, which also generate prediction distributions under the guidance of , resulting in two preference losses (denoted as and ). Moreover, a contrastive loss (denoted as ) is applied between the dynamic perception and preference activation paths, aiming to enhance the stability of the dual-path paradigm and improve the model’s discriminative ability by leveraging hard negative samples. During the inference stage, we sum the prediction distributions generated by the multimodal decoder from the temporal dynamic perception path and the audio/visual decoders from the global preference activation path, and then apply the argmax function to obtain the final answer.
@@ IV-A Experimental Setting
(a) Dataset and evaluation metric: In this paper, we conduct experiments on four representative AVQA benchmarks: MUSIC-AVQA [26], MUSIC-AVQA-R [36], MUSIC-AVQA-v2.0 [34], and AVQA [52], whose statistics are reported in Table I. MUSIC-AVQA is a widely used benchmark for audio-visual reasoning in music-performance scenes, containing 9,288 videos and 31,904/4,568/9,129 question-answer pairs for training, validation, and test, respectively. Since it is the most commonly adopted benchmark in prior work, we use it by default for our ablation and analysis experiments. MUSIC-AVQA-R is a robustness-oriented extension of MUSIC-AVQA, which keeps the same video set but expands the test split to 211,572 rephrased question-answer pairs, making it suitable for evaluating performance under more diverse question formulations and long-tail answer distributions. MUSIC-AVQA-v2.0 further improves the original benchmark by constructing a more balanced setting, and contains 10,492 videos with 37,408/5,346/10,819 question-answer pairs for training, validation, and test, respectively. We use it to assess the robustness of our method under reduced answer bias. To validate whether our model can generalize beyond music scenarios, we additionally evaluate on AVQA, a real-world benchmark containing 57,015 videos and 40,425/16,910 question-answer pairs for training/test. Following prior methods, we use answer accuracy (%) as the main evaluation metric across all datasets.
@@
Methods Params FLOPs Latency Acc (%) ST-AVQA 18.48M 3.19G 47.62ms 71.59 QA-Tiger 63.09M 9.57G 63.28ms 77.62 AV-Master 34.07M 6.20G 89.24ms 78.51
(b) Implementation details: For a fair comparison, we follow previous work [19] and adopt a similar setup: videos are uniformly sampled at a rate of 1 frame per second. Audio representations are extracted using the pre-trained VGGish model [17], while visual inputs and corresponding questions are encoded through the CLIP-ViT-L/14 model [39]. All extracted features are projected into a 768-dimensional space via a linear transformation. The predefined templates, and , have a dimension of 768. Both the visual and audio templates have a sequence length of 8 and are initialized using the visual and audio features at time step 0, respectively. These templates are then iteratively updated as time progresses within the AVFC module, where the time step corresponds to the length of the feature sequences. The model is optimized using Adam [20] and starts with a learning rate of 1e-4, which is reduced by a factor of 0.1 every 8 epochs. The batch size is set to 64, and the model is trained for 30 epochs. Our proposed model is trained on NVIDIA GeForce RTX 4090 and implemented in PyTorch.
@@ IV-B Quantitative Results
# Methods Average Accuracy (%) A-QA V-QA AV-QA Avg 1 w/o. ALL 73.37 79.23 69.82 72.94 2 w/o. AVFC 79.83 85.55 74.04 78.11 3 w/o. AVKF 79.76 85.84 74.06 78.19 4 w/o. DPCL 78.83 85.71 73.78 77.84 5 w/o. GPAP 77.90 85.05 73.10 77.12 6 w/o. TDPP 78.77 83.73 72.94 76.83 7 Dynamic Fus 77.53 84.97 73.78 77.41 8 AV-Master 79.95 86.58 74.22 78.51
@@
(c) MUSIC-AVQA-R: As shown in Table III, compared to the MUSIC-AVQA dataset, AV-Master demonstrates more significant improvements (+3.20%) on the MUSIC-AVQA-R dataset, achieving an overall accuracy of 71.19% and setting a new state-of-the-art performance. This significant performance gain can be attributed to the dual-path learning paradigm of AV-Master. Since MUSIC-AVQA-R substantially increases question diversity through large-scale rephrasing and further evaluates performance under answer head–tail imbalance, it favors models with stronger semantic grounding rather than those relying on fixed question templates. AV-Master benefits from this setting, as its TDPP builds answers upon temporally continuous evidence that is relevant to the question, while GPAP provides stable modality-level priors directly from raw unimodal features. As a result, our model is more robust to rephrased questions and distribution shifts, leading to greater improvements on MUSIC-AVQA-R.
@@
As shown in Table VI, AV-Master achieves competitive performance (78.5% accuracy) compared to various pretraining-based models, despite having significantly fewer parameters (34M) and not relying on large-scale pretraining. In contrast, methods such as CAT+ and AVicuna, which are based on large language models, achieve notably lower performance (50.1% and 49.6% accuracy) under zero-shot conditions. Although these approaches leverage abundant resources and powerful backbone models, they often exhibit weaker generalization capabilities in domain-specific audio-visual tasks. Other pretraining-based models achieve impressive performance (80.7% accuracy) after fine-tuning on the MUSIC-AVQA dataset. However, our method achieves comparable results using significantly less training data, fewer trainable parameters, and a more lightweight audio-visual feature encoder. This highlights the practicality of AV-Master and its suitability for resource-constrained audio-visual scenarios.
It also demonstrates that model size and the scale of pretraining data are not the only factors influencing AVQA performance. Although large pretrained models can learn broad cross-modal alignments, AVQA additionally requires fine-grained temporal localization and question-relevant modality selection. AV-Master introduces these task-specific inductive biases through the temporal dynamic perception path and the global preference activation path. As a result, even with only 34M parameters and without additional pretraining, AV-Master remains competitive with larger pretrained methods, highlighting the importance of architecture design for AVQA.
(g) Comparison of parameters and computational complexity: To further evaluate whether the performance improvement of AV-Master comes at the cost of efficiency, we compare it with representative methods in terms of parameter count, FLOPs, inference latency, and accuracy, as shown in Table VII. AV-Master achieves the best accuracy of 78.51% among the compared methods. Compared to ST-AVQA, AV-Master achieves a substantial performance gain (+6.92%) at the cost of acceptable overhead. We believe that this significant improvement more than justifies the marginal loss in efficiency. Compared with QA-TIGER, our method uses substantially fewer parameters (34.07M vs. 63.09M) and lower FLOPs (6.20G vs. 9.57G), while still achieving higher accuracy (78.51% vs. 77.62%). This suggests that the performance gain of AV-Master does not stem from simply scaling up the model size, but from the effectiveness of the proposed dual-path design.
On the other hand, the inference latency of AV-Master is higher than that of the compared baselines. We attribute this mainly to two factors: (1) the proposed temporal dynamic perception path performs iterative focus sampling over temporal segments, which introduces additional sequential computation. (2) the dual-path collaborative framework requires separate decoding and late fusion, which further increases runtime overhead. Therefore, although AV-Master is more favorable in terms of parameters and FLOPs, its wall-clock latency is relatively higher. Overall, these results demonstrate that AV-Master achieves a better trade-off between accuracy and model complexity, while future work may further optimize its inference efficiency through more parallelized implementation.